# Monotonic Paraphrasing Improves Generalization of Language Model Prompting

### v1

@@ 2.4 Search-based MonoPara
3 Experiment
3.1 Task I: Prompt Refinement
Task Description
To inspect if monotonic paraphrasing can refine a prompt and elicit better performance on prompting an LLM as designed, we use Super-NaturalInstruction (Sup-NatInst for short, Wang et al., 2022). This benchmark consists of 1,616 diverse NLP tasks and their expert-written prompts for evaluating the zero-shot generalizability of LLMs on a variety of NLP tasks. In Sup-NatInst, each task is paired up with an instruction that consists of the task definition for mapping an input text to a task output and several examples for demonstrating the desired or undesired output. To make our evaluation more challenging, we only use the task definition as a prompt. Since Sup-NatInst provides only one description of definition for each task, we use GPT4 Achiam et al. (2023) to generate $4$ more Sup-NatInst-like task descriptions for each involved task, which results in $5$ prompts for each task in total.
Datasets
Following Gonen et al. (2023), we choose $7$ classification tasks from Huggingface Datasets  [^2] https://huggingface.co/docs/datasets/index to have a set of diverse tasks: (i) GLUE Cola Warstadt et al. (2019) for grammatical acceptability discrimination; (ii) Newspop Moniz and Torgo (2018) for news classification; (iii)AG News Zhang et al. (2015) for news classification; (iv) IMDB Maas et al. (2011) for movie review sentiment analysis; (v) DBpedia Lehmann et al. (2015) for topic classification; (vi) Emotion Saravia et al. (2018) for tweet emotion classification; and (vii) Tweet Offensive Barbieri et al. (2020) for offensive tweet discrimination. We sample 1,000 random examples from each dataset for prompt evaluation. For these tasks, the prompt follows the input, and at the end of each prompt, we add the choices of classes following Gonen et al. (2023): “Choices: X, Y, Z. Answer: ” as it helps with accuracy. Defining the label space likely helps in the zero-shot setting because there are no previous demonstrations from which the model can learn the possible classes.
Evaluation Metrics
To inspect whether monotonic paraphrasing can refine a prompt into its low-perplexity counterpart and further enhance the generalization of LLM prompting, we use PPL (perplexity), Acc (accuracy), and BS (BERTScore, Zhang* et al., 2020) for evaluation. Perplexity. Following Gonen et al. (2023), we define the perplexity of the prompt as the perplexity of the full prompt sequence, including the input itself and the choices of labels. To avoid noise when computing perplexity, the prompts are instantiated with 1,000 random examples of the dataset (we keep the same selected examples for performance evaluation), and the perplexity is averaged across all instantiated prompts. BERTScore. As a paraphrasing scheme, MonoPara is also evaluated by its performance in semantic alignment with the source sentence. BERTScore computes a similarity score for each token in the paraphrase sentence with each token in the source sentence, which is evaluated based on pre-trained BERT Devlin et al. (2019) contextual embeddings. Accuracy. To compute accuracy, for each test sample, we obtain the model’s predicted probability of all classes and choose the highest ranking class as the prediction of the model. We use this metric over exact match for its stableness since the chances the model will generate an exact match of the label are relatively low and might appear in various formats. The accuracy is calculated over the 1,000 samples for each task.
Mistral 7B
@@ 3.1 Task I: Prompt Refinement
Starling 7B
Greedy BS 90.35 ± 0.67 88.35 ± 0.51 88.46 ± 0.98 90.04 ± 1.04 87.06 ± 0.79 90.38 ± 0.45 79.93 ± 11.52
Table 1: Results of Task I (Prompt Refinement). Ori. refers to the original prompts. Greedy refers to paraphrasing the original prompts without perplexity constraint. Mono-E refers to ensemble-based MonoPara. Mono-S refers to search-based MonoPara. The best accuracy for each task is in bold, while the lowest perplexity for each task is in blue.
Models
We study two auto-regressive models, Mistral-7B Jiang et al. (2023)and Starling-7B Zhu et al. (2023a), for their significant ability in instruction following and paraphrasing.
Results
As shown in Table 1, the two variants of MonoPara successfully refines prompts to those with perplexities lower than the original ones, thereby enhancing model performance on 10 out of 14 scenarios (2 models, each on 7 datasets). Specifically, Mono-E (ensembled-based MonoPara) improves the accuracy of Mistral-7B by an average of 2.64% and that of Starling-7B by an average of 1.85% on seven different datasets. In contrast, directly paraphrasing the prompts without perplexity constraint results in prompts with higher perplexities than the original ones in 13 out of 14 scenarios, most of which lead to worse task performance. This highlights the importance of incorporating perplexity constraint when paraphrasing prompts. Moreover, MonoPara achieves superior performance in terms of both average prompt perplexity and BertScore, indicating the synergistic role played by both the target model and the paraphrase model in prompt refinement. During decoding, the target model prioritizes achieving low perplexity, while the paraphrase model preserves the original prompt intent, resulting in a high BertScore.
3.2 Task II: Robustness Evaluation on Instruction Perturbation
Task Description
PromptBench introduces perturbation to task instructions of a diverse set of tasks, such as sentiment analysis, grammar correctness, duplicate sentence detection, and natural language inference. It includes perturbations at character-level using DeepWordBug Gao et al. (2018) to introduce typos, word-level using TextFooler Jin et al. (2020) to replace words with contextually similar words, and semantic-level by using prompts following the linguistic behavior of different languages.
Datasets
We select $4$ GLUE tasks Wang et al. (2018) from Huggingface Datasets for evaluation: (i) Cola, which is also used for Task I; (ii) SST-2 Socher et al. (2013) for two-way sentiment classification; (iii) MNLI Wang et al. (2018) for multi-genre inference identification; and (iv) MRPC Dolan and Brockett (2005) for paraphrase identification. Similar to the setting of Task I, the instruction follows the input, and at the end of each instruction, the choices of labels are appended: “Choices: X, Y, Z. Answer: .”
Model
For this task, we use Alpaca Taori et al. (2023), a model instruction-tuned from the LLaMA model Touvron et al. (2023a) on the 52k self-instruct dataset.
Results
As shown in Figure 3, we compare the performance of original prompts, prompts paraphrased without perplexity constraint, and our method Mono-E. Across all kinds of instruction perturbations, our method achieves consistent accuracy improvement by up to 10.13%. This demonstrates that MonoPara not only refines prompts for improved task performance but also enhances model robustness against various instruction perturbations.
3.3 Effect of Coefficient
We further analyze the effect of coefficient $\alpha$ between the target model and the paraphrase model in MonoPara. Table 2 presents a case study of $\alpha$ on the ensemble-based decoding variant (Mono-E) with a prompt under character-level perturbation. Obviously, a larger value of $\alpha$ amplifies the influence of the target model, resulting in a stronger perplexity constraint and monotonically lower perplexity. Consequently, the impact of the paraphrase model is diminished, leading to a lower BertScore. By adjusting $\alpha$, the user can find a balanced point to achieve better prompts that strike a balance between semantic fidelity to the original prompt intent and linguistic familiarity to the target model.
Content BertScore PPL
Ori Instruction
Read the provided excerpt and choose between
‘positive’ and ‘negative’ to describe its sentiment Iyk1lJt8yw:
- 78.12
Greedy Para Determine the sentiment of the given excerpt as either positive or negative 81.26 44.29
[…]
Figure 3: Model’s average accuracy across $4$ GLUE datasets, with each dataset having six instructions with perturbation added at character, word, and semantic levels. Mono-E has consistent improvement in accuracy across all types of perturbation compared to greedy paraphrasing.
4 Related Work
Different from existing paraphrasing techniques, this work conducts constrained decoding methods that seek to monotonically decrease the perplexity in the process of paraphrasing. This innovative approach serves as a general approach to rewrite and refine prompts or task instructions towards their more familiar counterparts for LMs, seeking to enhance the generalization of LM prompting without the need of any training effort.
Prompt Refinement. Prompt refinement has been actively explored in recent years with the aim of selecting, retrieving, or even generating prompts that lead to improved zero-shot performance of LM prompting. When an end-task objective is known, prior works on prompt refinement approaches often set the end-task performance as the objective, and leverage approaches such as gradient-based search, similar to continuous prompts but with projections onto a discrete vocabulary Shin et al. (2020). Other works have explored edit-based enumeration Prasad et al. (2023), reinforcement learning Deng et al. (2022), and large language model continuation and filtering Zhou et al. (2022). As for general purpose LLMs, a dedicated prompt refinement approach may not assume the pre-existance of any specific end tasks. Hence, more recent works Iter et al. (2023); Gonen et al. (2023); Wang et al. (2023); Lee et al. (2023) explored the heuristic criterion based on the familiarity of the language model with the language contained in the prompt, as measured by its perplexity. Lower perplexity prompts are preferred as they are expected to perform better across a wide range of tasks.
5 Conclusion
In this paper, we propose monotonic paraphrasing (MonoPara) for automatically and iteratively generating low-perplexity prompts for given LLMs, that can lead to better task performance. Following the high-level idea of paraphrasing prompts with perplexity constraints from the target LLM, we design ensemble-based and search-based decoding strategies for efficient prompt refinement. Experiments on prompt variations of unseen tasks and instructions demonstrate the effectiveness of MonoPara in reducing prompt perplexity while enhancing task performance and model robustness. Future work may apply MonoPara on other scenarios, such as searching stealthy prompts for red-teaming. Replacing perplexity with other criteria for different purposes of constrained generation are also meaningful research directions.
Limitations
The current investigation of MonoPara has the following limitations. First, using the target model as the paraphrase model is a natural technical choice. However, examining other paraphrasing models, especially those customized for paraphrasing, may lead to better performance. Second, we apply look-ahead decoding to improve search efficiency. However, exploring more advanced decoding methods, such as look-back decoding Xu et al. (2023), may also prove beneficial. Third, while we conduct experiments on a wide range of tasks, their original prompts and responses are relatively short. This actually limits the operational space of MonoPara. Experiments on tasks with longer inputs and outputs may provide additional evidence of MonoPara’s effectiveness.

### v3

@@ Appendix A Algorithm for Search-based MonoPara
Appendix B Results on Llama-3-8B
Results on Llama-3-8B are illustrated in Table 3.
Methods Metric GLUE Cola Newspop AG News Tweet Offensive
Ori. Acc 54.82 ± 1.73 73.82 ± 4.57 60.37 ± 7.59 63.25 ± 0.95
PPL 9.42 ± 5.08 17.72 ± 6.11 13.64 ± 4.97 23.64 ± 4.61
Para Acc 43.69 ± 9.02 68.45 ± 13.24 64.85 ± 3.73 64.22 ± 1.48
PPL 32.07 ± 8.43 29.41 ± 8.17 23.77 ± 4.32 46.14 ± 9.24
SPELL Acc 46.91 ± 2.33 71.67 ± 5.30 66.93 ± 3.70 64.73 ± 0.76
PPL 7.98 ± 1.58 12.47 ± 3.76 11.29 ± 2.79 8.79 ± 1.93
Mono-E Acc 58.35 ± 2.61 77.85 ± 4.01 65.45 ± 2.11 64.90 ± 0.59
PPL 6.64 ± 0.95 9.17 ± 2.59 8.29 ± 1.47 6.40 ± 3.61
Table 3: Performance of Llama-3-8B.
Appendix C Correlation Between Accuracy and Perplexity
To illustrate the interplay between task accuracy and prompt perplexity, we calculate the Pearson and Spearman correlation for each task based on the results in Table 1. The results are shown in Table 4. Most tasks show a strong negative correlation between perplexity and accuracy on both models except AG News, which is also discovered by Gonen et al. (2023) that for some of the tasks, there is not a negative correlation between perplexity and accuracy. The paraphrasing may be more helpful for cases where the nuances of expression don’t matter so much (or there’s enough room for error), but in domains where precision of language is important, perhaps perplexity-based paraphrasing might not be as helpful. AG News is a four-class news classification task with labels of “World”, “Business”, “Sci/Tech”, and “Sports”, which are not as easy as other binary classification tasks. As a result, the precision of the instruction for AG News is important. Additionally, we can see that the accuracy of AG News is not affected much under different methods, indicating that the accuracy is not so dependent on the perplexity of the prompt as other tasks. This indicates that MonoPara works best for tasks with high distinction among labels.
Model Metric GLUE Cola Newspop AG News IMDB DBpedia Emotion Tweet Offensive
Mistral Pearson -0.62 -0.88 0.19 -0.04 -0.70 -0.57 -0.48
Spearman -0.30 -0.99 0.30 -0.30 -0.39 -0.70 -0.30
Starling Pearson -0.83 -0.69 0.16 -0.60 -0.70 -0.96 -0.05
Spearman -0.90 -0.49 0.3 -0.3 -0.70 -0.89 -0.15
Table 4: Pearson and Spearman correlation between task accuracy and prompt perplexity w.r.t. the target model. Results are calculated based on Table 1.
Method Metric GLUE Cola Newspop AG News IMDB DBpedia Emotion Tweet Offensive
@@ Appendix C Correlation Between Accuracy and Perplexity
Para PPL 78.01 ± 11.22 49.37 ± 7.81 28.37 ± 2.58 22.27 ± 0.98 28.34 ± 1.36 46.53 ± 5.30 53.65 ± 1.92
Mono-E PPL 14.89 ± 2.75 8.33 ± 1.30 10.07 ± 2.93 13.05 ± 1.86 8.85 ± 0.88 13.67 ± 1.87 14.11 ± 3.53
Rerank (1)
Acc 62.20 ± 3.47 86.14 ± 3.14 69.31 ± 4.29 85.02 ± 4.13 69.88 ± 4.33 47.36 ± 3.28 66.12 ± 2.15
Budget = 10 PPL 14.27 ± 2.69 11.46 ± 1.99 11.92 ± 3.97 15.95 ± 2.83 13.15 ± 1.24 15.06 ± 2.71 15.63 ± 3.57
Acc 63.84 ± 2.46 85.86 ± 5.36 68.28 ± 5.16 86.52 ± 5.86 69.64 ± 3.53 47.60 ± 5.25 67.74 ± 1.45
Budget = 50 PPL 13.54 ± 3.16 10.53 ± 2.85 10.83 ± 3.24 13.44 ± 2.22 10.31 ± 1.85 14.14 ± 3.46 13.94 ± 3.86
Acc 64.58 ± 3.50 89.82 ± 2.04 72.33 ± 4.17 87.61 ± 4.93 72.52 ± 3.16 47.70 ± 3.56 68.5 ± 1.84
Budget = 100 PPL 10.22 ± 2.32 7.53 ± 1.82 9.52 ± 4.20 12.88 ± 3.01 8.16 ± 1.67 13.05 ± 1.34 11.56 ± 4.44
Rerank (2)
Acc 62.35 ± 4.77 86.96 ± 4.61 70.02 ± 5.27 84.86 ± 4.71 69.31 ± 3.94 46.10 ± 7.29 66.63 ± 4.53
Budget = 10 PPL 15.63 ± 3.85 12.27 ± 3.99 13.47 ± 3.63 16.43 ± 2.17 15.37 ± 2.14 16.64 ± 3.88 18.87 ± 4.48
Acc 62.08 ± 3.97 86.50 ± 5.10 71.12 ± 6.97 86.03 ± 5.24 70.13 ± 4.23 46.90 ± 5.83 65.15 ± 5.03
Budget = 50 PPL 14.42 ± 4.16 10.24 ± 4.36 11.69 ± 2.96 14.33 ± 3.55 13.62 ± 2.77 15.21 ± 3.38 18.53 ± 4.54
Acc 64.35 ± 4.91 87.82 ± 4.48 71.72 ± 6.07 86.64 ± 5.16 71.98 ± 3.85 47.18 ± 4.63 67.80 ± 8.95
Budget = 100 PPL 11.67 ± 4.37 8.93 ± 3.15 10.61 ± 4.89 13.47 ± 3.84 8.36 ± 3.57 14.41 ± 2.95 16.65 ± 3.53
Table 5: Reranking for ensemble-based MonoPara. The best results of each task are bolded and the second-best are italicized.
Newspop IMDB DBpedia Emotion Tweet
Ori 0.949 0.952 0.895 0.839 0.974
Para 0.813 0.951 0.835 0.645 0.871
Mono-E 0.956 0.895 0.927 0.931 0.968
Table 6: Pearson correlation between the perplexity of standalone prompt and its instantiated prompts. Pearson correlation is high (close to $1.0$) in most cases, indicating a high correlation.
Ori Para Mono-E
Un. 56.67 58.31 65.94
Cha. 51.03 (-9.95%) 53.62 (-8.04%) 65.81 (-0.20%)
Word 52.41 (-7.51%) 57.39 (-1.58%) 67.23 (+1.96%)
Sem. 49.42 (-12.79%) 54.22 (-7.01%) 64.35 (-2.4%)
Table 7: Performance degradation of Alpaca model compared with the unperturbed prompt. Un., Cha., and Sem. are short for unperturbed, character, and semantic, respectively, referring to different prompt perturbation methods.
Content BertScore
Ori Prompt In this task, you are presented with a short article. Your objective is to classify the article according to its category using the following labels: World, Sports, Business, Sci/Tech. Tag the text as “World” if it involves information about global events or issues. Mark it as “Sports” if it deals with sports-related content. Label it “Business” if it pertains to business affairs, markets, or economics. Assign “Sci/Tech” if it covers scientific or technological developments.
Para In this assignment, you are given a brief article. Your goal is to categorize the article based on its topic using these tags: World, Sports, Business, Sci/Tech. Identify it as “World” if it discusses international events or matters. Designate it as “Sports” if it focuses on athletic content. Categorize it as “Business” if it concerns commercial activities, markets, or the economy. Label it “Sci/Tech” if it reports on advancements in science or technology. 86.02
Mono-E In your task you will have an input which consists of a single article. You have to sort this article according to its category using a list which consists of 4 categories: “World”, “Sci/tech”, “Business”, or “Sports”. You should mark it with the appropriate label based on the content. The article should be labeled with the “Sports” category, when its topic deals with sports. When the article is about business, you need to mark it as a ’business’ label. If you are talking about science and tech developments then label the content with “Sci/Tech”, if it is a world-related issue or event then label it with “World”. 89.63
Table 8: Case study on the paraphrase results and BERTScore. Para tends to use new vocabulary for diversity, while Mono-E prefers words that are more familiar to the target model, which is often the same as in the original prompt.
Method Metric GLUE Cola Newspop AG News IMDB DBpedia Emotion Tweet
Ori. Acc 61.76 ± 2.93 85.28 ± 3.45 67.68 ± 3.47 84.12 ± 4.89 70.82 ± 2.27 44.32 ± 1.57 67.04 ± 0.95
PPL 15.72 ± 4.60 12.60 ± 2.40 12.74 ± 1.07 16.41 ± 0.80 8.10 ± 0.22 16.48 ± 2.14 19.25 ± 3.43
Para Acc 39.48 ± 8.19 76.72 ± 16.54 71.60 ± 2.70 86.60 ± 4.86 70.22 ± 5.71 43.65 ± 3.41 65.04 ± 1.46
@@
$\alpha$ = 0.1 Acc 60.80 ± 4.28 84.98 ± 3.41 69.09 ± 4.01 85.52 ± 1.72 70.42 ± 3.66 44.80 ± 2.31 65.56 ± 5.07
PPL 15.61 ± 2.51 12.32 ± 1.01 12.03 ± 2.96 15.70 ± 1.85 9.61 ± 0.61 16.16 ± 1.11 18.49 ± 2.84
$\alpha$ = 0.2 Acc 61.80 ± 5.63 85.32 ± 3.56 71.49 ± 4.32 85.08 ± 4.97 70.54 ± 5.23 44.04 ± 2.44 66.80 ± 5.19
PPL 15.22 ± 2.06 11.56 ± 1.27 11.68 ± 3.25 15.30 ± 1.55 9.23 ± 0.79 15.71 ± 0.20 16.92 ± 2.68
$\alpha$ = 0.3 Acc 61.54 ± 5.85 86.72 ± 4.94 72.4 ± 2.77 86.68 ± 4.21 71.90 ± 3.91 44.96 ± 2.67 67.56 ± 3.59
PPL 15.18 ± 2.84 10.21 ± 1.22 11.06 ± 2.74 14.42 ± 1.05 8.95 ± 0.75 15.21 ± 1.04 15.87 ± 3.61
$\alpha$ = 0.4 Acc 62.72 ± 4.37 89.44 ± 3.08 71.31 ± 2.27 87.86 ± 3.37 72.18 ± 3.44 45.58 ± 1.55 67.70 ± 1.60
PPL 14.93 ± 2.91 9.16 ± 0.57 10.39 ± 3.56 13.81 ± 2.60 9.04 ± 1.18 14.35 ± 0.90 15.18 ± 3.72
$\alpha$ = 0.6 Acc 64.30 ± 6.63 90.86 ± 3.24 71.66 ± 10.09 87.74 ± 4.04 71.36 ± 6.25 47.12 ± 1.66 66.26 ± 3.24
PPL 11.49 ± 1.00 7.53 ± 0.68 9.38 ± 0.88 12.15 ± 2.47 7.11 ± 1.08 13.12 ± 0.79 13.29 ± 2.58
$\alpha$ = 0.7 Acc 41.50 ± 13.68 54.52 ± 22.29 50.56 ± 21.95 64.84 ± 16.01 48.12 ± 16.26 38.62 ± 1.59 59.34 ± 7.75
PPL 8.80 ± 0.42 3.84 ± 0.91 3.43 ± 0.36 4.99 ± 0.28 2.34 ± 0.25 4.28 ± 0.55 6.58 ± 1.40
Table 9: Results with varying $\alpha$ values for Mono-E on Mistral-7B. The best accuracy for each task is in bold. For $\alpha$ = $0.4$ or $0.6$, the model shows outstanding performance on most of the datasets. But increasing alpha to $0.7$ results in a nonsensical prompt though the perplexity is super low, leading to corrupted task performance.
Appendix D Reranking for MonoPara
D.1 Implementation Setting
We implemented reranking on top of the linear combination of $P_{tar}$ and $P_{para}$ under two settings: (1) the coefficient $\alpha$ equals $0.5$, which is kept consistent with the results presented in Table 1 for direct comparison; (2) the coefficient is randomly selected from $[0.1,0.2,…,0.6]$ for a larger search space. In both cases, we allow sampling from the top $3$ tokens at every decoding step for randomness.
For each task, we use the $5$ prompts as the original prompts (the same setting as Table 1) and randomly generate $100$ paraphrases with the linear combination of $P_{tar}$ and $P_{para}$ for every single original prompt. For the generated $100$ paraphrases, we use the same scheme as is described in Section 3 to calculate the perplexity: each paraphrase is instantiated with $1,000$ random test samples, and the perplexity is averaged across all the $1,000$ instantiated prompts. For budget $k$, we select the first $k$ paraphrases from the $100$ candidates and choose the prompt with the lowest average instantiated perplexity to be tested on the task.
D.2 Analysis
For the reranking setting (1), with additional randomness when sampling paraphrases from the linear combination of $P_{tar}$ and $P_{para}$, a large enough budget (i.e., $100$ paraphrases) can consistently catch paraphrased prompts with lower perplexity which achieve better task performance (Table 5). However, with a small budget of $10$ or even $50$, reranking can hardly surpass Mono-E in either perplexity or task performance. The reason is that with randomness allowed, the paraphrase is no longer decoded greedily at each token, which undermines the constraint on the perplexity and leads to a large variation of the perplexity of the resulting paraphrase. This further affects the performance of these paraphrased prompts, which in turn proves the correlation between perplexity and task accuracy. Generally, we can expect vanilla Mono-E and reranking to break even with a large budget of around $50$ paraphrases.
The reranking setting (2) is not as effective as setting (1), where even a budget of $100$ could not surpass the vanilla Mono-E. The reason is that although a large search space is allowed with randomness on the combination coefficient $\alpha$, the constraint on perplexity while decoding is undermined even further. With more randomness for paraphrase sampling, the results of setting (2) also show a higher standard deviation than setting (1).
Appendix E Correlation Between Standalone and Instantiated Prompts
To bridge the intuition gap, we investigate the correlation between the standalone prompt and its instantiated prompts. We instantiate each task-specific prompt with $1,000$ input instances and the perplexity used for correlation evaluation is averaged over $5$ prompts for each task. As shown in Table 6, the Pearson correlation is high (close to 1.0) in most cases for different datasets despite slight noise for Emotion under greedy paraphrase of prompts. We can conclude that there is a high correlation between instantiated and standalone prompts. Thus, it is valid to use the perplexity of the standalone prompt optimized by $P_{tar}$ to serve as the constraint.
Appendix F Unperturbed Accuracy for Task II
Table 7 shows the task performance of three perturbation methods (Alpaca’s average accuracy across $4$ GLUE datasets) and performance degradation compared with the unperturbed prompt (as shown in the brackets). We can witness a narrow performance gap between perturbed and unperturbed original prompts for Mono-E, which is more robust than directly using the given prompt (Ori) or the greedily paraphrased one (Para). The reason is that Mono-E largely mitigates the effect of adversarial perturbations on the prompts and successfully rewrites the given prompts (whether perturbed or unperturbed) into higher-quality versions.
Appendix G BERTScore of Para and MonoPara
We provide a case study on the paraphrase results and the according BERTScore in Table 8. We could see that both Para and Mono-E largely preserve the semantic meaning of the original prompt. Further, Para tends to use new vocabulary for diversity, while Mono-E prefers words that are more familiar to the target model, which is often the same as in the original prompt. This might be the reason why Para results in lower BertScore compared to Mono-E and Mono-S.
Appendix H Ablation Study on Alpha Value
We provide results with various alpha values in Table 9, which illustrates the impact of alpha values on model accuracy and paraphrase perplexity across various datasets on the Mistral-7B model. With the increase of $\alpha$ value, the perplexity (PPL) of the paraphrase monotonically goes down due to the design and motivation of our method. Given the low perplexity, we compare the task accuracy of various alpha values. For $\alpha$ = $0.4$ or $0.6$, the model shows outstanding performance on some of the datasets. But increasing alpha to $0.7$ results in a nonsensical prompt though the perplexity is super low. From the results we can also see that the performance just corrupted with this nonsensical prompt. The setting of $\alpha$ should seek a balance between lower perplexity and the preservation of the original semantics. As a result, alpha = $0.5$ (Mono-E) offers a balanced and reliable performance, making it a suitable default choice.